# Fully Dynamic Matching and Ordered Ruzsa-Szemerédi Graphs

### v3

@@ 5.1 Linear Matchings
Take a matching $M\in\mathcal{M}^{\prime}$. Let $S_{M}$ be the set of pivots matched by $M$, i.e., $S_{M}=P\cap V(M)$. Let us define $Y_{M}:=V\setminus\cup_{v_{i}\in S}N_{i}$. From Observation 18, we get that all edges of $M$ must have both endpoints still in $Y_{M}$. Suppose for now that $M\in\mathcal{M}^{\prime}$ is chosen uniformly at random. We first argue that the expected size of $|Y_{M}|$ is relatively small (despite it preserving all edges of $M$ completely). We have
@@
$\leq n-\frac{4\delta c}{t}\cdot(tcn/2-n^{2}/x)$ (By 12.)
@@
Instead of the expected value of $Y_{M}$, we need some (weak) concentration. Take $\varepsilon=\delta c^{2}/2$ and note from 13 that $(1+\varepsilon)\operatorname{\textbf{E}}[|Y_{M}|]<(1+\varepsilon)(1-\delta c^{2})n<(1-\delta c^{2}/2)n$. We have
@@
We call a matching $M\in\mathcal{M}^{\prime}$ a good matching if $|Y_{M}|<(1-\delta c^{2}/2)n$. From 14 we get that at least $(\delta c^{2}/4)|\mathcal{M}^{\prime}|=\delta c^{2}t/8$ matchings are good.
@@
Now, we partition the good matchings in $\mathcal{M}^{\prime}$ into $2^{|P|}\leq 2^{x}$ classes depending on which subset of the pivots they match. Since there are at least $\delta c^{2}t/8$ good matchings, at least $\frac{\delta c^{2}t}{8\cdot 2^{x}}$ of them must belong to the same class $C$. Let $Y=Y_{M}$ for some $M\in C$. Note also that $Y=Y_{M^{\prime}}$ for any other matching $M^{\prime}\in C$ as well since they all match the same subset of pivots. Moreover, we have $|Y|<(1-\delta c^{2}/2)n$ since $M$ is good, and all matchings in $C$ only match vertices in $Y$ as discussed earlier (as a consequence of Observation 18). This implies that the graph $H$ on vertex set $Y$ and edge-set $C$ (i.e., including all edges of all matchings in $C$) is an $\mathsf{ORS}_{n^{\prime}}(cn,t^{\prime})$ graph for $n^{\prime}<(1-\delta c^{2}/2)n$ and $t^{\prime}\geq\delta c^{2}t/(8\cdot 2^{x})$ as desired. ∎
@@
$\geq\frac{c^{\prime\prime}n^{\prime}}{(1-\delta c^{\prime\prime 2}/2)}$ (By 16.)
@@
$\geq\Big{(}1+\delta\Big{(}(1-\delta c^{2}/10)c\Big{)}^{2}/2\Big{)}\Big{(}(1-\delta c^{2}/10)c\Big{)}n^{\prime},$ (Since the RHS is minimized when $c^{\prime\prime}$ is minimized, thus we can replace $c^{\prime\prime}$ with its LB from 15.)
$>(1+0.07c^{2})cn^{\prime}.$
We are now ready to prove Theorem 2.
Proof of Theorem 2.
$c_{k}$ $\geq(1+\kappa)^{k}c$
$=(1+\kappa)^{\log_{2}(1/c)/k}c$
$\geq 1.$
We prove by induction that for every $i\in[d]$,
$t_{i}\geq\frac{n}{\log^{(d-i)}_{b}n}.$
$t_{i}$ $\geq t_{i-1}/2^{0.5\beta(n_{i-1}/c_{i-1}t_{i-1})}$
$=t_{i-1}/b^{0.5(n/t_{i-1})}$ (Since we defined $b=2^{\beta/c}$.)
$=\frac{n}{\log^{(d-i+1)}_{b}n}\cdot b^{-0.5\log^{(d-i+1)}_{b}n}$
$=\frac{n}{\log^{(d-i+1)}_{b}n\cdot\sqrt{\log^{(d-i)}_{b}n}}$
$\geq\frac{n}{\log^{(d-i)}_{b}n},$
concluding the proof.
5.2 When Matchings are (Very) Large
The goal of this section is to prove that the number of matching of an $\mathsf{ORS}$ graph with $r>n/4$ is bounded by a constant value. More precisely we prove Theorem 4. An equivalent of this theorem was already known for RS graphs (see [29]), and we show that almost the same proof carries over to ORS graphs as well.
Theorem 3.
For any $c>1/4+\varepsilon$, $\mathsf{ORS}_{n}(cn,t)\leq 1/\varepsilon+1$.
First, we argue that the number of common vertices of any two matching in ORS is at most $r$, using this we can prove that a constant value bounds the number of matching since the number of subsets of length $2r$ with at most $r$ common vertices between any two sets.
Lemma 16.
Suppose $G$ is an $\mathsf{ORS}(r,t)$ graph on $n$ vertices and let $M_{1},\ldots,M_{t}$ be the corresponding ordered list of matchings as in Definition 2. Then for any $i\not=j$ in $[t]$, we have $|V(M_{i})\cap V(M_{j})|\leq r$.
Proof.
Let us assume w.l.o.g. that $i<j$ and suppose that $|V(M_{i})\cap V(M_{j})|\geq r+1$. By the pigeonhole principle, this means that there must be an edge in $M_{i}$ whose both endpoints are matched in $M_{j}$, contradicting the fact from Definition 2 that $M_{j}$ is an induced matching in $M_{1}\cup\ldots\cup M_{j}$. ∎
Now, we use a similar approach as [29] to prove that $t<1/\varepsilon$. This theorem only uses the fact that the number of common vertices of two matchings is at most $r$.
Proof of Theorem 4.
Using Lemma 16 we have that the distance between sets of vertices of any two matchings is at most $r$, this implies a constant upper bound for the size of $\mathsf{ORS}$ using bounds from coding theory. To improve the upper bound we set $v_{i}\in\{0,1\}^{n}$, a vector that represents vertices in $V(M_{i})$. Let us use $a_{i}$ to denote the number of vectors $v_{j}$ with 00 in their $i$-th coordinate and let $b_{i}$ be the number of vectors with $1$ in their $i$-th coordinate. If we set $v_{0}$ a vector of all zeros. Note that we have $a_{i}+b_{i}=t+1$. Now, by double counting we have:
$2r{t+1\choose 2}\leq\sum_{0\leq i<j\leq t}dist(v_{i},v_{j})=\sum_{i=1}^{n}a_{i}b_{i}.$ (10)
By $dist(v_{i},v_{j})$ we mean the Hamming distance which satisfies the following inequality:
$dist(v_{i},v_{j})=|V_{i}|+|V_{j}|-2|V_{i}\cap V_{j}|\geq 2r.$
Note that in 10, $a_{i}b_{i}$ maximizes when we have $a_{i}=b_{i}=(t+1)/2$ for odd $t$ and $a_{i}=(t+2)/2,b_{i}=t/2$ for even $t$. This concludes that $t<1/\varepsilon+1$
6 Bounding Density of (Ordered) Ruzsa-Szemerédi Graphs
6.1 Linear Matchings
Lemma 17.
Observation 18.
Proof.
We are now ready to prove Lemma 17.
Proof of Lemma 17.
We explain the proof in a few steps.
Step 1: The pivots.
$|P|\leq x.$ (11)
$|E_{0}|-\sum_{i=1}^{k}|N_{i}|^{2}\leq n^{2}/x.$
$\sum_{i=1}^{k}|N_{i}|^{2}\geq|E_{0}|-n^{2}/x\geq tcn/2-n^{2}/x.$ (12)
Step 2: Vertex reduction.
Take a matching $M\in\mathcal{M}^{\prime}$. Let $S_{M}$ be the set of pivots matched by $M$, i.e., $S_{M}=P\cap V(M)$. Let us define $Y_{M}:=V\setminus\cup_{v_{i}\in S}N_{i}$. From Observation 18, we get that all edges of $M$ must have both endpoints still in $Y_{M}$. Suppose for now that $M\in\mathcal{M}^{\prime}$ is chosen uniformly at random. We first argue that the expected size of $|Y_{M}|$ is relatively small (despite it preserving all edges of $M$ completely). We have
$\leq n-\frac{4\delta c}{t}\cdot(tcn/2-n^{2}/x)$ (By 12.)
$=n-2\delta c^{2}n+\frac{4\delta cn^{2}}{tx}$
$=n-2\delta c^{2}n+\delta c^{2}n$
$=(1-\delta c^{2})n.$ (13)
Instead of the expected value of $Y_{M}$, we need some (weak) concentration. Take $\varepsilon=\delta c^{2}/2$ and note from 13 that $(1+\varepsilon)\operatorname{\textbf{E}}[|Y_{M}|]<(1+\varepsilon)(1-\delta c^{2})n<(1-\delta c^{2}/2)n$. We have
$\geq 1-\frac{1}{1+\varepsilon}$ (By Markov’s inequality.)
$\geq\varepsilon/2=\delta c^{2}/4.$ (14)
We call a matching $M\in\mathcal{M}^{\prime}$ a good matching if $|Y_{M}|<(1-\delta c^{2}/2)n$. From 14 we get that at least $(\delta c^{2}/4)|\mathcal{M}^{\prime}|=\delta c^{2}t/8$ matchings are good.
Step 3: Grouping good matchings.
Now, we partition the good matchings in $\mathcal{M}^{\prime}$ into $2^{|P|}\leq 2^{x}$ classes depending on which subset of the pivots they match. Since there are at least $\delta c^{2}t/8$ good matchings, at least $\frac{\delta c^{2}t}{8\cdot 2^{x}}$ of them must belong to the same class $C$. Let $Y=Y_{M}$ for some $M\in C$. Note also that $Y=Y_{M^{\prime}}$ for any other matching $M^{\prime}\in C$ as well since they all match the same subset of pivots. Moreover, we have $|Y|<(1-\delta c^{2}/2)n$ since $M$ is good, and all matchings in $C$ only match vertices in $Y$ as discussed earlier (as a consequence of Observation 18). This implies that the graph $H$ on vertex set $Y$ and edge-set $C$ (i.e., including all edges of all matchings in $C$) is an $\mathsf{ORS}_{n^{\prime}}(cn,t^{\prime})$ graph for $n^{\prime}<(1-\delta c^{2}/2)n$ and $t^{\prime}\geq\delta c^{2}t/(8\cdot 2^{x})$ as desired. ∎
Lemma 19.
Proof.
Case 1: $\alpha\geq\delta c^{3}/10$.
and
Case 2 – $\alpha<\delta c^{3}/10$:
$n^{\prime}$ $<(1-\delta c^{\prime\prime 2}/2)n^{\prime\prime},$ (16)
the number of matchings satisfies
and finally since Lemma 17 does not change the size of matchings, we get
$c^{\prime}n^{\prime}$ $=c^{\prime\prime}n^{\prime\prime}$
$\geq\frac{c^{\prime\prime}n^{\prime}}{(1-\delta c^{\prime\prime 2}/2)}$ (By 16.)
$\geq(1+\delta c^{\prime\prime 2}/2)c^{\prime\prime}n^{\prime}$
$\geq\Big{(}1+\delta\Big{(}(1-\delta c^{2}/10)c\Big{)}^{2}/2\Big{)}\Big{(}(1-\delta c^{2}/10)c\Big{)}n^{\prime},$ (Since the RHS is minimized when $c^{\prime\prime}$ is minimized, thus we can replace $c^{\prime\prime}$ with its LB from 15.)
@@ 6.2 When Matchings are (Very) Large
For any $c>1/4+\varepsilon$, $\mathsf{ORS}_{n}(cn,t)\leq 1/\varepsilon+1$.

### v4

@@ 5.1 Linear Matchings
Take a matching $M\in\mathcal{M}^{\prime}$. Let $S_{M}$ be the set of pivots matched by $M$, i.e., $S_{M}=P\cap V(M)$. Let us define $Y_{M}:=V\setminus\cup_{v_{i}\in S}N_{i}$. From Observation 14, we get that all edges of $M$ must have both endpoints still in $Y_{M}$. Suppose for now that $M\in\mathcal{M}^{\prime}$ is chosen uniformly at random. We first argue that the expected size of $|Y_{M}|$ is relatively small (despite it preserving all edges of $M$ completely). We have
@@
$\leq n-\frac{4\delta c}{t}\cdot(tcn/2-n^{2}/x)$ (By 5.)
@@
Instead of the expected value of $Y_{M}$, we need some (weak) concentration. Take $\varepsilon=\delta c^{2}/2$ and note from 6 that $(1+\varepsilon)\operatorname{\textbf{E}}[|Y_{M}|]<(1+\varepsilon)(1-\delta c^{2})n<(1-\delta c^{2}/2)n$. We have
@@
We call a matching $M\in\mathcal{M}^{\prime}$ a good matching if $|Y_{M}|<(1-\delta c^{2}/2)n$. From 7 we get that at least $(\delta c^{2}/4)|\mathcal{M}^{\prime}|=\delta c^{2}t/8$ matchings are good.
@@
Now, we partition the good matchings in $\mathcal{M}^{\prime}$ into $2^{|P|}\leq 2^{x}$ classes depending on which subset of the pivots they match. Since there are at least $\delta c^{2}t/8$ good matchings, at least $\frac{\delta c^{2}t}{8\cdot 2^{x}}$ of them must belong to the same class $C$. Let $Y=Y_{M}$ for some $M\in C$. Note also that $Y=Y_{M^{\prime}}$ for any other matching $M^{\prime}\in C$ as well since they all match the same subset of pivots. Moreover, we have $|Y|<(1-\delta c^{2}/2)n$ since $M$ is good, and all matchings in $C$ only match vertices in $Y$ as discussed earlier (as a consequence of Observation 14). This implies that the graph $H$ on vertex set $Y$ and edge-set $C$ (i.e., including all edges of all matchings in $C$) is an $\mathsf{ORS}_{n^{\prime}}(cn,t^{\prime})$ graph for $n^{\prime}<(1-\delta c^{2}/2)n$ and $t^{\prime}\geq\delta c^{2}t/(8\cdot 2^{x})$ as desired. ∎
@@
$\geq\frac{c^{\prime\prime}n^{\prime}}{(1-\delta c^{\prime\prime 2}/2)}$ (By 9.)
@@
$\geq\Big{(}1+\delta\Big{(}(1-\delta c^{2}/10)c\Big{)}^{2}/2\Big{)}\Big{(}(1-\delta c^{2}/10)c\Big{)}n^{\prime},$ (Since the RHS is minimized when $c^{\prime\prime}$ is minimized, thus we can replace $c^{\prime\prime}$ with its LB from 8.)
@@ 5.2 When Matchings are (Very) Large
For any $c>1/4+\varepsilon$, $\mathsf{ORS}_{n}(cn)\leq 1/\varepsilon+1$.